Show that X^2 + Y^2 + Dx + Ey + F=0 represents a circle of positive radius if and only if D^2 + E^2 - 4F > 0.

1
Expert's answer
2022-04-12T12:35:46-0400

Show that "x^2+y^2+Dx+Ey+F=0" represents a circle of positive radius if and only if "D^2 + E^2 - 4F > 0" .

Solution:

The equation of the circle can be written as

"(x-a)^2+(y-b)^2=R^2" , where "(a, b)" is a centre of the circle, R is a radius of the circle.

Let's transform our equation to the form of the equation of the circle:

"x^2+y^2+Dx+Ey+F=x^2+2\frac D2x+\frac{D^2}{4}+""y^2+2\frac E2y+\frac{E^2}{4}-\frac{D^2}{4}-\frac{E^2}{4}+F=0"

"(x+\frac D2)^2+(y+\frac E2)^2-\frac{D^2}{4}-\frac{E^2}{4}+F=0"

"(x+\frac D2)^2+(y+\frac E2)^2=\frac{D^2}{4}+\frac{E^2}{4}-F" (*)

Compare (*) and the equation of the circle. Equation (*) represents a circle if and only if "\frac{D^2}{4}+\frac{E^2}{4}-F=R^2" . "R^2" is always a positive value so we can conclude

"\frac{D^2}{4}+\frac{E^2}{4}-F>0"

or

"D^2+E^2-4F>0"